a man is 20 years older than his son. In ten years, He will be twice as old as his son will then be. How old is each?
Answer by mananth(16946)

a man is 20 years older than his son. In ten years, He will be twice as old as his son will then be. How old is each?
son = x years
father = x+20
...
after 10 years
son = x+10
father = x+20+10 = x+30 years
...
x+30=2(x+10)
x+30=2x+20
2x-x=30-20
x= 10 son's age
father's age = x+20 = 10+20 = 30 years
...
m.ananth@hotmail.ca